# Is Less More? Exploring Token Condensation as Training-free Adaptation for CLIP

### v2

@@ 3.2 Token Condensation as Adaptation
Domain-aware token reservoir $\mathfrak{R}$. Notably, both $f_{\operatorname{prune}}$ and $f_{\operatorname{merge}}$ rely on a class-specific token reservoir =$\mathfrak{R}=\{\mathfrak{R}_{c}\}_{c=1}^{C}$. Each buffer $\mathfrak{R}_{c}=\{(\mathbf{H}_{c}(\mathbf{z}_{i};\mathbf{t}_{c}),\mathbf{A}_{i,c}^{\operatorname{cls}})\}_{i=1}^{M}$ is structured as a priority queue that retains the top $M$ most reliable domain <cls> tokens of target samples, which serve to implicitly distil semantic information from the corresponding text prompt $\mathbf{t}_{c}$ to guide the visual adaptation. These domain <cls> tokens are crucial alignment proxies: although the architectures of the text encoder $E_{t}$ and the visual encoder $E_{v}$ differ, the selected domain <cls> tokens help determine which visual tokens best align with text features from CLIP’s perspective. The reliability of these domain <cls> tokens is quantified by entropy scores $\mathbf{H}_{c}(\mathbf{z}_{t},\mathbf{t}_{c})$,
@@
which act as keys to update the reservoir $\mathfrak{R}_{c}$. At each time step $t$, for each visual embedding $\mathbf{z}_{t}$, the corresponding <cls> embeddings from all $L$ layers $\mathbf{A}_{t,c}^{\operatorname{cls}}=[\mathbf{v}_{\operatorname{cls}}^{1},\ldots,\mathbf{v}_{\operatorname{cls}}^{L}]\in\mathbb{R}^{L\times D_{v}}$ will be stored in $\mathfrak{R}_{c}$ if $\operatorname{argmax}(\mathbf{p}_{t,c})=c$, ensuring that only the most semantically consistent samples are retained:
@@
Logits Self-correction. To counter the shifts on the semantic side, we introduce a logits self-correction mechanism that leverages domain <cls> tokens stored in $\mathfrak{R}$. In particular, the <cls> embedding of the current sample, $\mathbf{V}_{t}^{\operatorname{cls}}\in\mathbb{R}^{L\times D_{v}}$, is compared with the stored domain <cls> tokens denoted as a set $\mathcal{A}=\{\mathbf{A}_{i,c}^{\operatorname{cls}}\}_{i=1}^{M}$. The cosine similarity between these cross-layer <cls> tokens serves as a token-level classifier, which provides auxiliary information to adjust the predicted probability $\mathbf{p}_{t,c}$ from a visual perspective:
@@
where $\lambda$ is the logit correction weight. $\mathbbm{1}_{c}\in\mathbb{R}^{C}$ the one-hot vector for the $c$-the class. The layer-specific exponential scaling coefficients are denoted as $\mathbf{P}=[\exp(\frac{l}{\beta})]_{l=1}^{L}\in\mathbb{R}^{L}$, where $\beta$ controls the influence of different layers. We show that this correction temperature $\beta$ provides semantic interpretability, as further discussed in Sec. 4.3. This self-correction mechanism ensures better alignment between the visual and semantic contexts, improving robustness in handling semantic shifts.
Figure 3: An overview of domain-aware cross-head token pruning.
3.3 Domain-aware Cross-head Token Pruning
Prior token pruning methods for ViTs such as [26] primarily discard patch tokens with lower averaged attention scores $\mathbf{S}\in\mathbb{R}^{N}$ relative to the <cls> token $\mathbf{v}_{\operatorname{cls}}^{l}$ across all attention heads. The pruning process can be expressed as
$\hat{\mathbf{V}}_{\operatorname{prune}}^{l}\leftarrow\{\hat{\mathbf{v}}_{i}^{l}~{}|~{}\mathbf{S}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\in[N]\},$ (7)
$~{}\mathbf{S}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{Attention}(\mathbf{v}_{\operatorname{cls}}^{l}\mathbf{W}_{Q}^{h},\mathbf{v}^{l}_{i}\mathbf{W}_{K}^{h}),$
where $\hat{\mathbf{V}}_{\operatorname{prune}}^{l}$ denotes the set of tokens retained after pruning at layer $l$. Here, the pruning threshold $\theta_{\operatorname{prune}}(\alpha,R)$ is determined by the desired pruning ratio, ensuring only top-ranked $\alpha\cdot R\cdot N$ tokens are retained. However, this approach faces two limitations when applied in TTA tasks: (1) The <cls> token is universal and may not be specifically aligned with the target class set. It may capture broad, unrelated semantics (e.g., “cat food”), leading to the retention of irrelevant tokens that mislead the model into making incorrect predictions of the target class (e.g., “cat”). (2) Averaging attention scores across all heads risks omitting important details, as each attention head tends to focus on distinct features (e.g., shape, color). Outliers in attention heads (highlighted by red circles in Fig. 3) may disproportionately dominate the overall score, overshadowing valuable information by other heads and leading to suboptimal pruning decisions. To overcome these limitations, we propose a cross-head token pruning function that evaluates the token importance individually for each attention head and utilizes the averaged relative ranking positions to determine which tokens to prune (see Fig. 3). This approach reaches a more robust cross-head consensus and mitigates the impact of outliers. To guide pruning with more representative <cls> tokens, we sample an domain <cls> token for the $(l-1)$-th layer from $\mathfrak{R}_{c^{*}}$, where $c^{*}$ is determined by comparing the cosine similarity between the current $\mathbf{v}^{l}_{\operatorname{cls}}$ token embedding and the stored domain <cls> tokens $\mathbf{A}^{l-1}_{i,c}\in\mathcal{A}$:
$\mathbf{A}_{c^{*}}^{l-1}=\frac{1}{M}\sum\nolimits_{i\in[M]}\mathbf{A}_{i,c^{*}}^{l-1},$ (8)
where $\mathbf{A}_{c^{*}}^{l-1}\in\mathbb{R}^{D_{v}}$ is the averaged domain <cls> token sampled from $\mathfrak{R}_{c^{*}}$, providing historical context that is better aligned with target semantics. $c^{*}=\arg\max\nolimits_{c\in[C]}\operatorname{cos}(\mathbf{v}_{\operatorname{cls}}^{l},\mathbf{A}_{i,c}^{l-1}),$. Subsequently, we refine the attention map by inserting this historical domain <cls> token $\mathbf{A}_{c^{*}}^{l-1}$ to compute the pruning scores $\mathbf{S}^{\operatorname{head}}_{i}$ for the $i$-th token:
$\mathbf{S}^{\operatorname{head}}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{rank}_{h}(i),$ (9)
$~{}\hat{\mathbf{V}}_{\operatorname{prune}}^{l}\leftarrow\{\hat{\mathbf{v}}_{i}^{l}~{}|~{}\mathbf{S}^{\operatorname{head}}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\in[N]\},$
where $\text{rank}_{h}(i)$ gives the relative ranking position of token $i$ in head $h$ based on the attention score to the concatenation of $\mathbf{v}^{l}_{\operatorname{cls}}$ and $\mathbf{A}_{c^{*}}^{l-1}$, computed as $\operatorname{Attention}([\mathbf{v}_{\operatorname{cls}}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{Q},[\mathbf{V}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{K})$. [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] indicates concatenation. This method ensures that tokens receiving consistently high attention across individual heads are retained, thereby achieving greater robustness to outliers in specific attention heads.
3.4 Domain-aware Token Merging
As depicted in Fig. 2(a), a subset of tokens, although relevant to the target class, exhibit high uncertainty. These tokens are referred to as class-ambiguous tokens, identified from the ranked token list derived using LABEL:eq:sort:
where $\theta_{\operatorname{merge}}(R)$ denotes thresholds for token selection during merging. The selected tokens $\mathbf{V}^{l}_{\Phi}=\{\mathbf{v}_{i}^{l}\}_{i\in\Phi}$ can introduce variance or noise into latent representation $\mathbf{z}_{t}$ and negatively impact the final classification decision. To address this, we propose a domain-aware token merging strategy to consolidate these tokens into more representative ones.
Existing token merging strategies either truncate neighbored token pairs with high similarity with bipartite soft matching [3] or apply spectral clustering [2], and graph pooling [58] to merge similar tokens at higher costs. In contrast, we adopt a more efficient coreset selection approach, which identifies the most representative tokens $\hat{\mathbf{V}}^{l}_{\operatorname{merge}}\in\mathbb{R}^{K\times D_{v}}$ from $\mathbf{V}^{l}_{\Phi}$ and assigns the remaining ambiguous tokens to these selected tokens. The coreset selection strategy is equivalent to solving the K-Center problem [57, 43]. The objective is to select $K$ center tokens such that the maximum distance between any token and its nearest center is minimized. Formally, the greedy search for optimization is defined as follows:
where $\mathcal{N}(k)$ represents the set of tokens assigned to center $k$. The value of $K$ is kept small, with $K\ll N$, allowing our merging algorithm to operate with linear complexity. The final embedding is composed of the class token $\mathbf{v}^{l}_{\operatorname{cls}}$, the pruned tokens $\hat{\mathbf{V}}^{l}_{\operatorname{prune}}$ excluding those in the class-ambiguous set $\Phi$, and the merged tokens $\hat{\mathbf{V}}^{l}_{\operatorname{merge}}$. This combined embedding is then passed to the next layer in the ViT. The study of $k$ and the details of the algorithm are provided in the supplementary material.
@@ 4.1 Experimental Setup
Datasets. Following prior works, we conducted the cross-dataset (CD) benchmark. The CD benchmark assesses the model’s performance on unseen classes across 10 datasets: Aircraft [30], Caltech101 [10], Cars [23], DTD [6], EuroSAT [15], Flower102 [33], Food101 [4], Pets [35], SUN397 [60], and UCF101 [46]. Details of other benchmark experiments are provided in the supplementary material.
@@
Implementation Details. We utilize the official CLIP prompts as text inputs. The batch size is set to 1 without data augmentations to mimic realistic deployment scenarios. All experiments are conducted using the pre-trained CLIP models, specifically using ViT-B/16 and ViT-L/14 architectures as the visual backbone. We set $K$ to 2. Notably, our method is training-free, which achieves rapid adaptation with minimized hyperparameter dependency. All experiments are performed on one NVIDIA RTX A6000 GPU.
4.2 Main Results
Table 1: Results on the cross-dataset benchmark using CLIP ViT-B/16, including the number of learnable parameters (L-Param.) for learning-based TTA methods. denotes the averaged GFLOPs across all datasets.
@@ 4.2 Main Results
Tab. 1 presents the results for fine-grained cross-dataset benchmark using the ViT-B/16 architecture. As observed in Fig. 2(a), the core idea behind TCA is that condensing inattentive tokens can effectively mitigate distribution shifts caused by visual-text misalignment. This concept is first validated by the improved performance of token pruning baselines over CLIP inference, where a condensed token set yields a $0.9\%$ increase in average accuracy when $R=0.9$. TCA further enhances its performance by dealing with visual-text misalignment, moving visual features toward historical domain <cls> tokens from DTR. As a result, TCA achieves an average accuracy of 68.69%, outperforming both train-required and training-free baselines without augmentation. Conventional TTA methods perform poorly on all datasets even with the requirement of fine-tuning a large amount of learnable parameters. In contrast, prompt-tuning methods, although requiring fewer learnable parameters, rely heavily on augmentation and struggle to effectively handle visual shifts. While TDA is a training-free method, it requires a large number of hyperparameters (a total of 10 for managing positive and negative caches) to achieve optimal performance. On the other hand, TCA uses significantly fewer hyperparameters and delivers a 1.72% improvement in average accuracy over TDA, with approximately 12.2% fewer GFLOPs. Further details on TDA combined with token condensation baselines, as well as the impact of the visual backbone (ViT-L/14) and OOD benchmarks, are presented in supplementary material.
@@ 4.3 Ablation Study
We conducted a comprehensive ablation study to evaluate TCA’s effectiveness and efficiency. For further analysis on merging center and pruning to merging ratio, see the supplementary material.
Table 2: Impact of correction weight $\lambda$.
$\lambda$ 2 3 4 5 6 7 8
Pets 89.53 89.32 89.13 88.96 88.96 88.66 88.44
Flower102 72.43 72.76 73.20 73.16 73.16 73.33 73.16
EuroSAT 60.15 65.74 68.80 69.51 69.84 70.16 70.43
Table 3: Impact of scale factor $\beta$.
$\beta$ 0.01 0.05 1 3 5
Pets 89.51 89.53 89.37 89.42 89.26
Flower102 73.33 73.08 70.93 70.56 70.44
EuroSAT 63.64 64.06 69.86 70.26 70.43
Impact of Logits Correction Temperature $\beta$. Tab. 3 examines how different logit correction temperatures $\beta$ affect the adaptation results. The intuition is that with a smaller $\beta$ value, the logits correction will emphasize the tokens in shallower layers (Eq. 6), while a larger $\beta$ value will shift the focus to deeper layers. We observe that a smaller value of $\beta$ is preferred for the Pets dataset as it contains animals as objects, requiring more high-level contextual information for accurate predictions [37]. In contrast, for EuroSAT, the best predictions are obtained with larger $\beta$ values, suggesting that low-level, local information is crucial. This aligns well with the nature of the dataset, where different types of land can be distinguished by features such as colors and edges. Nevertheless, our method consistently provides significant improvements across all $\beta$ values, with accuracy gains of up to 20%, highlighting the effectiveness of logit correction using the domain <cls> tokens.
@@
Impact of GFLOPs Budget. We evaluate TCA’s performance under different GFLOPs budgets: $R=\{0.6,0.7,0.8,0.9\}$, resulting in GFLOPs of 9.91, 11.68, 13.27, and 15.45, respectively, compared to the baseline ($R=1$, 17.58 GFLOPs). As shown in Fig. 4(b), condensing inattentive tokens can even enhance performance on certain datasets, notably Pets, and EuroSAT. Specifically for EuroSAT, when $R=0.9$, the model’s adaptation performance is significantly improved, aligning with our findings in Fig. 2(a). However, excessively aggressive pruning budgets (e.g., GFLOPs less than 13) lead to significant performance degradation across all datasets. This occurs since higher pruning rates may inadvertently remove informative tokens, causing irreversible harm in training-free scenarios where we lack supervision or the ability to update the model for extra correction.
Table 4: Impact of reservoir updating strategy.
@@
Table 5: Impact of the proposed components in TCA.
@@
Impact of Component. The impact of the historical domain <cls> $\mathbf{A}_{c^{*}}^{l}$ and the head-wise sorting score $\mathbf{S}^{\text{head}}$ (Eq. 8) is presented in Tab. 5. We observe that each component individually contributes to performance improvements. On the Food101 and Pets datasets, incorporating either component yields measurable gains in accuracy. By leveraging historical domain <cls> tokens, the model acquires rich contextual information, enhancing the stability of token importance over time. Simultaneously, cross-head token sorting ensures that token pruning decisions are more robust by accounting for consensus across attention heads. An intriguing case arises with the EuroSAT dataset. Here, the baseline performance without any components is 68.14%. Applying either component alone results in a slight performance decrease. However, when both components are used together, performance significantly improves to 70.43%. This outcome emphasizes the necessity of combining historical domain <cls> tokens and cross-head token sorting to fully realize the model’s potential.

### v3

@@ 3.3 Domain-aware Token Reservoir
To keep representative anchor tokens at domain level, we introduce domain-aware token reservoir =$\mathfrak{R}=\{\mathfrak{R}_{c}\}_{c=1}^{C}$. Notably, both token reduction and logits correction are building on the top of these saved tokens. In $\mathfrak{R}$, each buffer $\mathfrak{R}_{c}=\{(\mathbf{H}_{c}(\mathbf{z}_{i};\mathbf{t}_{c}),\mathbf{A}_{i,c}^{\operatorname{cls}})\}_{i=1}^{M}$ is structured as a priority queue that retains the top $M$ most reliable domain anchor tokens of target samples, which serve to implicitly distil semantic information from the corresponding text prompt $\mathbf{t}_{c}$ to guide the visual adaptation. These domain anchor tokens are crucial alignment proxies: although the architectures of the text encoder $E_{t}$ and the visual encoder $E_{v}$ differ, the selected domain anchor tokens help determine which visual tokens best align with text features. The reliability of these domain anchor tokens is quantified by entropy scores,
@@
which act as keys to update the reservoir $\mathfrak{R}_{c}$. At each time step $t$, for each visual embedding $\mathbf{z}_{t}$, the corresponding anchor embeddings from all $L$ layers $\mathbf{A}_{t,c}^{\operatorname{cls}}=[\mathbf{v}_{\operatorname{cls}}^{1},\ldots,\mathbf{v}_{\operatorname{cls}}^{L}]\in\mathbb{R}^{L\times D_{v}}$ will be stored in $\mathfrak{R}_{c}$ if $\operatorname{argmax}(\mathbf{p}_{t,c})=c$, ensuring that only the most semantically consistent samples are retained:
@@
3.4 Domain-aware Cross-head Token Reduction
Inspired by our preliminary studies as despite in Fig. 2(a), we introduce token condensation for training-free adaptation. Prior token reduction methods [27] primarily discard patch tokens with lower averaged attention scores $\mathbf{S}\in\mathbb{R}^{N}$ relative to the <cls> token $\mathbf{v}_{\operatorname{cls}}^{l}$ across all attention heads $\mathbf{S}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{Attention}(\mathbf{v}_{\operatorname{cls}}^{l}\mathbf{W}_{Q}^{h},\mathbf{v}^{l}_{i}\mathbf{W}_{K}^{h})$. However, this approach faces two limitations when applied in TTA tasks: (1) The <cls> token is universal and may not be specifically aligned with the target class set. It may capture broad, unrelated semantics (e.g., “cat food”), leading to the retention of irrelevant tokens that mislead the model into making incorrect predictions of the target class (e.g., “cat”). (2) Averaging attention scores across all heads risks omitting important details, as each attention head tends to focus on distinct features (e.g., shape, color). Outliers in attention heads (highlighted by red circles in Fig. 3) may disproportionately dominate the overall score, overshadowing valuable information by other heads. To overcome these limitations, we propose a domain-aware cross-head token reduction that evaluates the token importance individually for each attention head with the consideration of saved domain anchor tokens and utilizes the averaged relative ranking positions to determine which tokens to prune and which to merge(see Fig. 3). This approach reaches a more robust cross-head consensus and mitigates the impact of outliers.
Domain-aware Token Evaluation. Firstly, we sample the domain anchor token for the $(l-1)$-th layer from domain-aware class buffer $\mathfrak{R}_{c^{*}}$, where $c^{*}$ is determined by the maximum cosine similarity between the current $\mathbf{v}^{l}_{\operatorname{cls}}$ token embedding and the average stored domain anchor tokens in $\mathfrak{R}_{c}$ for each class $c$. Here, the averaged domain anchor token is calculated by $\mathbf{A}_{c}^{l-1}=\frac{1}{M}\sum\nolimits_{i\in[M]}\mathbf{A}_{i,c}^{l-1}$. $c^{*}=\arg\max\nolimits_{c\in[C]}\operatorname{cos}(\mathbf{v}_{\operatorname{cls}}^{l},\mathbf{A}_{i,c}^{l-1})$. Subsequently, we refine the attention map by concatenating $\mathbf{v}^{l}_{\operatorname{cls}}$ with the historical domain anchor token $\mathbf{A}_{c^{*}}^{l-1}$,
$\operatorname{Attention}([\mathbf{v}_{\operatorname{cls}}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{Q},[\mathbf{V}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{K}),$ (5)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] indicates concatenation. This providing historical context that is better aligned with target semantics.
Figure 3: An overview of domain-aware cross-head token pruning.
Cross-head Token Reduction. To better evaluate the attentiveness of the tokens, we compute the token reduction score $\mathbf{S}^{\operatorname{head}}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{rank}_{h}(i)$ for the $i$-th token, where $\text{rank}_{h}(i)$ gives the relative ranking position of token $i$ in head $h$ based on the attention score in Eq. 5. This ensures that tokens receiving consistently high attention across individual heads are retained, thereby achieving greater robustness to outliers in specific attention heads.
We find out the class-irrelevant and class-ambiguous tokens by the cross-head score $\mathbf{S}^{\operatorname{head}}_{i}$ and formulate token reduction in TCA as:
where $f_{\operatorname{prune}}(\cdot;\mathfrak{R}):\mathbb{R}^{N+1}\mapsto\mathbb{R}^{(\alpha\cdot R\cdot N)+1}$ and $f_{\operatorname{merge}}(\cdot;\mathfrak{R}):\mathbb{R}^{(\alpha\cdot R\cdot N)+1}\mapsto\mathbb{R}^{R\cdot N+1}$ are pruning and merging respectively, responsible for reducing the number of tokens from $N+1$ (including the <cls> token) to $R\cdot N+1$, where $R$ is the fraction of tokens to be preserved. $\alpha$ controls the extent of token pruning. We first prune the class-irrelevant ones:
$\vspace{-1ex}\hat{\mathbf{V}}_{\operatorname{prune}}^{l}\leftarrow\{\hat{\mathbf{v}}_{i}^{l}~{}|~{}\mathbf{S}^{\operatorname{head}}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\in[N]\},$ (7)
where $\hat{\mathbf{V}}_{\operatorname{prune}}^{l}$ is the set of tokens retained after pruning at layer $l$. The pruning threshold $\theta_{\operatorname{prune}}(\alpha,R)$ allows only top-ranked $\alpha\cdot R\cdot N$ tokens are retained.
As depicted in Fig. 2(a), the class-ambiguous tokens, although relevant to the target class, exhibit high uncertainty:
where $\theta_{\operatorname{merge}}(R)$ denotes thresholds for token selection during merging. The selected tokens $\mathbf{V}^{l}_{\Phi}=\{\mathbf{v}_{i}^{l}\}_{i\in\Phi}$ can introduce variance or noise into latent representation $\mathbf{z}_{t}$ and negatively impact the final classification decision. To address this, we propose a domain-aware token merging strategy to consolidate these tokens into more representative ones. Here, rather than truncating neighbored token pairs with bipartite soft matching [3], applying spectral clustering [2], or graph pooling [63], we adopt a more efficient coreset selection approach. Details of the coreset selection process and algorithm are provided in the supplementary material.
3.5 Logits Self-correction
@@
To counter the shifts on the semantic side after token condensation, we introduce a logits self-correction mechanism that leverages domain anchor tokens stored in $\mathfrak{R}$. In particular, the visual <cls> token of the current sample (if using CLIP) $\mathbf{V}_{t}^{\operatorname{cls}}\in\mathbb{R}^{L\times D_{v}}$ is compared with the stored domain anchor tokens $\mathcal{A}=\{\mathbf{A}_{i,c}^{\operatorname{cls}}\}_{i=1}^{M}$. The cosine similarity between these cross-layer tokens serves as a token-level classifier, which provides auxiliary information to adjust the predicted probability $\mathbf{p}_{t,c}$ from a visual perspective:
@@
where $\lambda$ is the logits correction weight. $\mathbbm{1}_{c}\in\mathbb{R}^{C}$ the one-hot vector for the $c$-the class. The layer-specific exponential scaling coefficients are denoted as $\mathbf{P}=[\exp(\frac{l}{\beta})]_{l=1}^{L}\in\mathbb{R}^{L}$, where $\beta$ controls the influence of different layers. We show that this correction temperature $\beta$ provides semantic interpretability, as further discussed in Sec. 4.3. This self-correction mechanism ensures better alignment between the visual and semantic contexts, improving robustness in handling semantic shifts.
@@ 4.1 Experimental Setup
Datasets. Following prior works, we evaluate our method on the cross-dataset (CD) benchmark, which measures model performance on unseen classes across 10 datasets: Aircraft [32], Caltech101 [10], Cars [24], DTD [6], EuroSAT [15], Flower102 [36], Food101 [4], Pets [38], SUN397 [65], and UCF101 [49]. Additionally, we assess TCA’s robustness to distribution shifts using CIFAR-100-Corrupted [16] (CIFAR-100-C), which introduces varying corruption severities. Further details on additional experiments are provided in the supplementary material.
@@
Implementation Details. For CLIP, we use its official prompts. The batch size is set to 1, without data augmentations, to mimic realistic deployment scenarios. All experiments are conducted using pre-trained CLIP models with ViT-B/16 and ViT-L/14 architectures as the visual backbone. We set $K$ to 2. For SigLIP and SigLIP v2, we select models with ViT-B/16 as the backbone. Since the SigLIP series learns visual embeddings differently (i.e., via MAP rather than relying on the last layer’s <cls> token), we adapt our method accordingly. Specifically, we use pooled attention weights to obtain the domain anchor and the token attentiveness score ($S_{i}^{\text{head}}$) for cross-head token reduction and logits self-correction. Unless otherwise stated, all ablation studies are conducted on CLIP with ViT-B/16. Notably, our method is training-free, enabling rapid adaptation with minimal hyperparameter dependency. All experiments are performed on a single NVIDIA RTX A6000 GPU.
Table 1: Results on the cross-dataset benchmark using CLIP ViT-B/16, including the number of learnable parameters (L-Param.) for learning-based TTA methods. denotes the averaged GFLOPs across all datasets. The best performance (aug-free) is bolded.
@@
4.2 Main Results
Tab. 1 presents the results for fine-grained cross-dataset benchmark using the ViT-B/16 architecture on CLIP. As observed in Fig. 2(a), the core idea behind TCA is that condensing inattentive tokens can effectively mitigate distribution shifts caused by visual-text misalignment. This concept is first validated by the improved performance of token pruning baselines over CLIP inference, where a condensed token set yields a $0.9\%$ increase in average accuracy when $R=0.9$. TCA further enhances its performance by dealing with visual-text misalignment, moving visual features toward historical domain anchor tokens from DTR. As a result, TCA achieves an average accuracy of 68.69%, outperforming both train-required and training-free baselines without augmentation. Conventional TTA methods perform poorly on all datasets even with the requirement of fine-tuning a large amount of learnable parameters. In contrast, prompt-tuning methods, although requiring fewer learnable parameters, rely heavily on augmentation and struggle to effectively handle visual shifts. While TDA is a training-free method, it requires a large number of hyperparameters (a total of 10 for managing positive and negative caches) to achieve optimal performance. On the other hand, TCA uses significantly fewer hyperparameters and delivers a 1.72% improvement in average accuracy over TDA, with approximately 12.2% fewer GFLOPs. Further details on the impact of the visual backbone (ViT-L/14) and OOD benchmarks are provided in the supplementary material.
@@ 4.3 Ablation Study
We conducted a comprehensive ablation study to evaluate TCA’s effectiveness and efficiency. For further analysis on hyperparameter, see the supplementary material.
Results on Various Severity Levels. To evaluate the effectiveness of TCA across different shift levels, we conduct experiments on five severity levels of three corruption types in the CIFAR-100-Corrupted dataset. Our results in Tab. 2 reveal a performance decline for EViT across most severity levels, particularly for snow and brightness corruptions. This may be due to the CLIP model embedding perturbed samples differently when token reduction is applied, leading to misinterpretations. In contrast, our proposed TCA consistently demonstrates improvements, highlighting the effectiveness of the DTR module in preserving informative and reliable token representations for logits correction.
Table 2: Improvements over CLIP inference on CIFAR-100-C.
[…]
$\operatorname{Severity}$ $\operatorname{CLIP}$ $\operatorname{EViT}$ $\operatorname{Ours}$ $\operatorname{CLIP}$ $\operatorname{EViT}$ $\operatorname{Ours}$ $\operatorname{CLIP}$ $\operatorname{EViT}$ $\operatorname{Ours}$
1 31.90 -3.17% +2.16% 35.34 -2.49% +1.50% 41.00 -2.10% +2.10%
2 20.67 +0.68% +3.58% 29.72 -1.99% +1.14% 41.44 -1.54% +2.85%
3 15.05 -1.53% +7.04% 29.14 -1.68% +1.61% 41.83 -1.82% +2.37%
4 8.85 -0.34% +11.41% 27.04 -2.51% +1.48% 41.12 -1.82% +4.74%
5 2.69 +1.49% +18.59% 24.85 -0.80% +3.30% 38.10 -2.13% +4.75%
Results on Multimodal VLMs. To evaluate the generalizability of our proposed method, we conduct additional experiments on SigLIP and SigLIP v2. As shown in Tab. 3, TCA consistently improves the performance of both SigLIP and SigLIP v2 without requiring model tuning. Notably, EuroSAT achieves a significant gain of 21.12% over SigLIP direct inference, aligning with trends observed in Tab. 1. On average, TCA yields improvements of 2.21% for SigLIP and 1.32% for SigLIP v2 across datasets, demonstrating its potential. While minor degradations appear in Food101, overall performance remains stable or improved across most datasets, reinforcing TCA’s robustness.
Table 3: Improvements in Cross-Dataset benchmark over SigLIP and SigLIP v2 inference on ViT-B/16.
Method Aircraft Caltech101 Cars DTD EuroSAT Flower102 Food101 Pets SUN397 UCF101 Average
SigLIP 40.50 97.44 90.71 62.83 39.86 84.13 89.06 93.10 69.65 70.84 73.81
+TCAR=0.9 +0.74% +0.58% +0.12% +2.44% +21.12% +1.39% -0.17% +0.56% +2.02% +3.40% +2.21%
SigLIP v2 50.05 97.61 93.36 62.41 42.09 85.30 90.14 94.88 73.34 72.09 76.13
+TCAR=0.9 +0.42% +0.25% +0.11% +2.76% +11.17% +0.39% -0.04% +0.02% +0.64% +3.22% +1.32%
@@
Impact of GFLOPs Budget. We evaluate TCA under different GFLOPs budgets: $R=\{0.6,0.7,0.8,0.9\}$, resulting in GFLOPs of 9.91, 11.68, 13.27, and 15.45, respectively, compared to the baseline ($R=1$, 17.58 GFLOPs). As shown in Fig. 4(b), condensing inattentive tokens can even enhance performance on certain datasets, notably Pets, and EuroSAT. Specifically for EuroSAT, when $R=0.9$, the model’s adaptation performance is significantly improved, aligning with our findings in Fig. 2(a). However, excessively aggressive pruning budgets (e.g., GFLOPs less than 13) lead to significant performance degradation across all datasets. This occurs since higher pruning rates may inadvertently remove informative tokens, causing irreversible harm in training-free scenarios where we lack the ability to update the model for further correction.
Table 4: Impact of reservoir updating strategy with CLIP.
@@
Table 5: Ablation study of the proposed components in TCA.
@@
Impact of Component. The impact of the domain anchor tokens $\mathbf{A}_{c^{*}}^{l}$ and the head-wise sorting score $\mathbf{S}^{\text{head}}$ (Sec. 3.4) is presented in Tab. 5. We observe that each component individually contributes to performance improvements. On the Food101 and Pets datasets, incorporating either component yields measurable gains in accuracy. By leveraging historical domain anchor tokens, the model acquires rich contextual information, enhancing the stability of token importance over time. Simultaneously, cross-head token sorting ensures that token pruning decisions are more robust by accounting for consensus across attention heads. An intriguing case arises with the EuroSAT dataset. Here, the baseline performance without any components is 68.14%. Applying either component alone results in a slight performance decrease. However, when both components are used together, performance significantly improves to 70.43%. This outcome emphasizes the necessity of combining historical domain anchor tokens and cross-head token sorting to fully realize the model’s potential.
Comparison with TDA $+$ Token Condensation. We evaluate the performance for TDA $+$ token pruning and merging baselines and show the performance gain over TDA $+$ ATS in Fig. 5. Although TDA achieves considerable performance gain, it heavily relies on the negative cache and a large set of hyperparameters. In contrast, TCA’s accuracy gain significantly surpasses that of TDA $+$ EViT and TDA $+$ ToME across multiple datasets and on average, even with a minimal set of hyperparameters, highlighting its superior adaptation capability.